OpenStudy (anonymous):

The total rainfall in Forks, Washington, this October was 8 3/8 in. In September, 60% less rain fell than in October. How much rain fell in September in Forks, Washington? Express your answer as a mixed number in simplest form.

OpenStudy (anonymous):

>> the question does not make sense

OpenStudy (anonymous):

now it should

undeadknight26 (undeadknight26):

3/8 = 3.75 3.75 * .225?

OpenStudy (anonymous):

I need it as a fraction

OpenStudy (anonymous):

lol u can make it to a fraction now its easy

undeadknight26 (undeadknight26):

You got the rest of this @hhmh4 <3

OpenStudy (anonymous):

kk

OpenStudy (wmj259):

8-3/8 inches? * .60% =5.025

undeadknight26 (undeadknight26):

Ohhhhh 8 (3/8) I missed that part ._. Nice catch @wmj259 !

OpenStudy (anonymous):

5 1/40 tht wht i got

OpenStudy (wmj259):

.028 = 1/40 as a fraction so just combine that to 5 to get 5-1/40 inches

OpenStudy (wmj259):

@undeadknight26 thank you!